OpenStudy (anonymous):

What is the sum of this geometric series? 1/3-1/9+1/27-1/81+...

OpenStudy (dumbcow):

formula for sum of geometric series \[S = \frac{a(1-r^{n})}{1-r}\] a = 1st term r = ratio n = num of terms

OpenStudy (dumbcow):

in this case n is infinity , i assume "..." means it goes on forever

OpenStudy (dumbcow):

a = 1/3 r = -1/3